OpenStudy (anonymous):

Find the particular solution of the equation which y = 1 when x = 0, and then give this particular solution in explicit form.

OpenStudy (anonymous):

\[36y= (\exp ^{(-3x+2sinx)/6})^{6}\]

OpenStudy (unklerhaukus):

\[36y=\left( e^{\frac{-3x+2\sin x}{6}}\right) ^6=e^{{-3x+2\sin x}}\]

OpenStudy (anonymous):

Yes.

OpenStudy (unklerhaukus):

a good first first i would think

OpenStudy (unklerhaukus):

step*

OpenStudy (anonymous):

Can you run through how you got that result?

OpenStudy (unklerhaukus):

\[\left(e^{\frac{-3x +2\sin x}{6}}\right)^6\]\[=\left(e^{\frac{-3x +2\sin x}{6}}\right)\left(e^{\frac{-3x +2\sin x}{6}}\right)\left(e^{\frac{-3x +2\sin x}{6}}\right)\left(e^{\frac{-3x +2\sin x}{6}}\right)\left(e^{\frac{-3x +2\sin x}{6}}\right)\left(e^{\frac{-3x +2\sin x}{6}}\right)\]\[=\left(e^{6\left(\frac{-3x +2\sin x}{6}\right)}\right)\]\[=\left(e^{-3x +2\sin x}\right)\]

OpenStudy (anonymous):

It is already integrated.

OpenStudy (unklerhaukus):

i dont really understand the question

OpenStudy (unklerhaukus):

shouldn't there be a constant /

OpenStudy (anonymous):

Ohhhh! Yes it should be +C in the right hand side within bracket. Sorry!

OpenStudy (anonymous):

\[36y= (\exp ^{(-3x+2sinx)/6} +C)^{6}\]

OpenStudy (unklerhaukus):

how did i know that? because we are solving for c such that y = 1 when x = 0 this was impossible before as 36≠1]

OpenStudy (unklerhaukus):

\[36y=\left(e^\frac{(−3x+2 \sin x)}{6}+C\right)^6\]

OpenStudy (anonymous):

Yes

OpenStudy (unklerhaukus):

when y=1, x=0 \[36y=\left(e^\frac{(-3(0)-2sin0)}{6}+C\right)^6\]\[36=\left( e^0+C \right) ^6\]\[36=\left( 1+C \right) ^6\]

OpenStudy (unklerhaukus):

\[\sqrt[6]{36}=1+C\]

OpenStudy (unklerhaukus):

\[C=\sqrt[6]{36}-1\]\[=36^{1/6}-1\]

OpenStudy (anonymous):

Great.

OpenStudy (unklerhaukus):

\[=(6^2)^{1/6}\]\[=6^{1/3}=\sqrt[3]{6}\]

OpenStudy (unklerhaukus):

so \[C=\sqrt[3]{6}-1\]

OpenStudy (anonymous):

3rd root. Can see very well

OpenStudy (anonymous):

Can't see very well small print. I assume it is

OpenStudy (anonymous):

many thanks to you.

OpenStudy (unklerhaukus):

\[{\large C=\sqrt[3]{6}-1=6^{\left(\frac 1 3\right)} -1}\]

OpenStudy (unklerhaukus):

\[\frac{dy}{dx}= (2 \cos x − 3) e^{(−3x+2 \sin x)/6} y^{5/6}?\]

OpenStudy (anonymous):

Yes

OpenStudy (unklerhaukus):

what is the y^5/6 thing doing?

OpenStudy (anonymous):

*y^5/6 after expo

OpenStudy (unklerhaukus):

it is just timesed on is it?

OpenStudy (anonymous):

part 1 (2cos) . part 2 (exp). part3 (y^)

OpenStudy (anonymous):

Hope that is clear

OpenStudy (unklerhaukus):

it is clear now